OpenStudy (anonymous):

simplify 2uv^8 x 5x^7 x 3v^2 x^-6 u^-5

OpenStudy (anonymous):

To indicate multiplication, use an asterisk (*) or otherwise use equation format.

OpenStudy (anonymous):

http://prntscr.com/1vjf68

OpenStudy (anonymous):

30*u^-4*v^10*x

OpenStudy (anonymous):

just multiply the numbers then multiply the variables a^m * a^n = a ^ (m+n)

OpenStudy (anonymous):

says to use positive exponents only

OpenStudy (anonymous):

a^- n = 1/a^n

OpenStudy (anonymous):

so, 30 * v^10 * x / u^4

OpenStudy (anonymous):

whats *

OpenStudy (anonymous):

30 x v^10 x x/u^4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Nvm

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

* = multiply hahha